LEP EXPERIMENTS COMMITTEE
Minutes of the 46th meeting, 9 September 1997

OPEN SESSION

1. Status of LEP2 and present performance

M. Lamont reported on the status of LEP2. The fire in the SPS surface building resulted in a 35 day delay in LEP startup, with about 20 days of physics lost in 1997 (79 days c.f. 98 days foreseen). The machine is presently operated at a beam energy of 91.5 GeV with 90/60 optics and 4 bunches per beam. So far a total of 15.5 pb-1 has been delivered at high energy, with good overall performance: initial luminosities exceeding 4.0x1031 cm-2s-1 and a maximum integrated luminosity of 1.4 pb-1 per day. However the beam current is still moderate; the peak value for physics has been 4.8 mA. The search for a good low emittance lattice continues during MD periods - both by improving the 108/90 optics and by developing a new 102/90 optics, for which the initial studies are encouraging. Concerning the LEP2 energy calibration programme, good (24%) polarisation has been achieved at 45 GeV but only limited (5%) polarisation at 50 GeV.

Some vacuum chamber problems have been experienced with the increased synchrotron radiation power at high energies, resulting in a deterioration of beam lifetimes in collision to around 5 h. A systematic programme of reconditioning vacuum pumps and installing additional instrumentation and cooling is underway. OPAL has experienced severe off-momentum backgrounds and DELPHI high synchrotron radiation doses to the silicon vertex detector.

2. Status of the LEP2 RF system

G. Geschonke reported that the sc RF system is progressing very well in performance and reliability. A total of 240 sc and 86 copper cavities are presently installed in LEP, providing a total accelerating voltage of 2.6 GV. With the presently installed cryogenic cooling power of 6.6 kW, the maximum total current is 6.8 mA for 4 bunches per beam, or 9.6 mA for 8 bunches. A further 32 sc cavities have been delivered to CERN and are planned for installation in the 97/98 shutdown. By the start of the 1999 cycle a further 16 sc cavities should have been installed, reaching the final total of 288 sc cavities. About 98% of these cavities reach above the nominal field of 6 MV/m.

3. Estimates of LEP2 future performance

D. Brandt reported on the estimates of LEP2 performance in 1998 and beyond. Assuming 90/60 optics and allowing for 2 klystrons to be off (and 10% power safety margin), the expected energy in 1998 is sqrts = 189 GeV. The performance in 1998 is expected to be limited by the available cryogenic power (6.2 kW), resulting in a peak luminosity of 8x1031 cm-2s-1 and an integrated luminosity of between 1.0 pb-1 per day (at an efficiency eta=0.15) and 1.4 pb-1 per day (eta=0.20). The performance in 1999 will depend on the optics and installed cryogenic power. To achieve an integrated luminosity of 150 pb-1 will require a peak luminosity of (7-10)x1031 cm-2s-1, depending on the efficiency. This seems achievable. To reach an energy near sqrts = 200 GeV will require a cryogenic power of 12 kW and operating the Nb-film sc RF cavities near 7 MV/m.

4. LEP working group reports

               Electroweak physics:  R. Clare

                Higgs particles:  P. Igo-Kemenes

                Four-jet events:  D. Schlatter

Reports were presented from three of the LEP working groups. A wealth of combined electroweak results have been prepared which now include LEP2 as well as LEP1 data (http://www.cern.ch/LEPEWWG). Within the Standard Model, these data indicate an upper limit on the higgs mass, mH < 420 GeV (95% CL). The direct lower limit from the combined 4 LEP experiments is mH > 77 GeV (95% CL) - a limit that is about 7 GeV higher than any single measurement. The anomalous 4-jet signal seen only by ALEPH remains a puzzle. It has been demonstrated that the other 3 detectors have efficiencies and sensitivities for these events comparable to ALEPH's. Statistically it is equally unlikely (7x10-4 probability) that they are due to a fluctuation in standard processes or that they are due to a new physics signal.

5. ALEPH request for 130/136 GeV running

P. Dornan presented a request from ALEPH to run at 130/136 GeV in order to resolve the 4-jet puzzle. This involves an integrated luminosity of 12 pb-1 divided between the energies 130 and 136 GeV, which corresponds to an additional factor of two in luminosity compared with the earlier data. The preferred running time is before the end of 1997, and certainly before LEP2 has reached its maximum energy. These energies are chosen since they correspond to the strongest and clearest observation of a signal by ALEPH (9 events with a background of 1 event). Combining the 4 detectors and based on the previous 130/136 GeV data, a 12 pb-1 sample will yield an estimated 25 signal events on top of a background of 7 events.

CLOSED SESSION

Present: J. Dainton, J. Drees, L. Foà, G. Goggi, K. Hübner, P.O. Hulth, G. Kantardjian,
              J. Kirkby (Secretary), K.-H. Kissler, W. Lohmann, T. Lohse, M. Mangano, J. May,
              J. Panman, L. Pape, Y. Sirois, I. Videau, P. Wells and P. Zerwas (Chairman).

Apologies: S. Bethke, F. Gasparini, R. Marshall, T. Müller, S. Myers, S. Pokorski, M. Turala
              and C.H. Llewellyn Smith.

1. Approval of the minutes of the 45th meeting

The minutes of the 45th meeting (LEPC 97-4/LEPC 45) were approved without modification.

2. Chairman's report

The Chairman thanked the departing members of the committee - M. Calvetti, P. Hansen, M. Klein and B. Mansoulié - and welcomed the new members: F. Gasparini, P.O. Hulth, T. Müller and Y. Sirois. He also thanked S.Bethke, J. Panman and I. Videau for having accepted to continue on the committee for an additional year. He congratulated the machine teams on the rapid cleanup and commissioning of LEP after the fire in the SPS surface building.

The Chairman and the Research Director reported on the current situation regarding LEP operation in 2000. Both the Research Board and the SPC agree with the conclusion of the LEPC that there is a very significant discovery potential in the Higgs and SUSY sectors by raising the energy of LEP near to 200 GeV and collecting about 200 pb-1 per experiment at this energy. Since the peak energy will not be reached until the start of the 1999 cycle, the Research Board and the SPC have supported operation of LEP in 2000, subject to securing the required funds. Special contributions have been requested for operating LEP in 2000, and a final decision is expected in December 1997.

3. Report from the LEP Coordinator

The LEP Coordinator reported that LEP had achieved a rapid startup on 18 July after the fire. The 1997 schedule is now expected to comprise about 85 days for physics and 14 days for machine development. After providing 2.5 pb-1 of Z calibration data per experiment, a total of 15.5 pb-1 had been delivered in 25 days at sqrts = 183 GeV. Assuming LEP continues at this energy and performance (0.6 pb-1 per day) for the rest of this year, the high energy data could total about 50 pb-1, and it is reasonable to hope the performance will improve. The estimated performance at 130/136 GeV is about 1.0 pb-1 per day, with about 2 days setup required to make the transition from high energy operation. This would imply about 2 weeks to collect the 12 pb-1 requested by ALEPH.

In 1998 LEP is expected to operate for 157 days in the period from 11 May until 1 November, comprising about 10 days for 2.5 pb-1 of Z calibration data, about 117 days for high energy running (sqrts = 189 GeV) and about 30 days for machine development. Based on the 90/60 optics, the integrated luminosity at high energy is estimated to be between 1 and 1.4 pb-1 per day.

Concerning the beam energy calibration at LEP2, the precision in 1996 at sqrts = 172 GeV is sigma(Ebeam) = 30 MeV. This is based on the flux loop/NMR measurements of the LEP dipoles, which are precisely calibrated vs. energy at lower beam energies by resonant depolarisation measurements. These have been made at 45 GeV and 50 GeV, and indicate a relative error of 2 x 10-4. Assuming a linear extrapolation to higher energies, this indicates an error of 15 MeV at 86 GeV. The present estimated error has been inflated by a factor of 2 to account for uncertainties in the reproducibility of the measurements and in the assumption of a linear extrapolation. With the same assumptions, and before further measurements and improvements are made, the precision at sqrts = 183 GeV is sigma(Ebeam) = 35 MeV. Further resonant depolarisation measurements are underway - both at 45 and 50 GeV as before and also at lower (40 GeV) and higher (55 GeV) energies - in order to improve the lever-arm and investigate the linearity. Depending on the outcome of these measurements, the error could be reduced to sigma(Ebeam) = 18 MeV.

The LEP Energy Working Group is also investigating independent methods that may be used to cross-check and improve the resonant depolarization/flux loop/NMR measurement of the beam energy at LEP2. One technique is to measure the radiative returns to the Z which results in an expected error of 30-40 MeV from the combined four detectors. Another approach under study is a precision spectrometer measurement of the beam energy using a dipole in the LEP ring. The latter could achieve a precision of 10 MeV assuming the beam position could be measured to 1 um with precision beam position monitors (which would be calibrated with Z data) and assuming the magnetic field of the dipole can be measured to 10-4 precision at energies between 45 and 100 GeV.

4. Discussion on the open session reports

Concerning the machine, it was noted that the performance from 1998 onwards will be limited by cryogenic cooling power for the sc RF cavities. Assuming the upgrade to 12 kW cooling power is completed prior to the 1999 startup then LEP could reach an energy near 200 GeV that year.

The committee congratulated the LEP Working Groups on their results and look forward to hearing the reports from the QCD and SUSY working groups at the next meeting in November.

5. Discussion on the ALEPH energy request

Concerning the ALEPH request for 12 pb-1 data at 130/136 GeV, the committee recognises the importance of resolving the 4-jet puzzle and that further data at these low energies are required to determine whether these events are due to a fluctuation of standard processes or to new physics. Furthermore the committee agrees on the importance of settling this question as soon as possible, also on the grounds that future LEP data will be taken at higher energies which have increased discovery potential. Questions were raised whether a lower integrated luminosity would be sufficient to resolve the puzzle, and it was agreed that this could be answered by the new data themselves, according to pre-defined criteria.

The committee therefore made the following recommendation for how to proceed. A meeting should be held during the MD period at the start of October (week 40) including the LEP Spokesmen, the Chairman of the 4-jet Working Group, the Directors of Research and Accelerators, the LEP Coordinator, and the Chairman and Secretary of the LEPC. The purpose would be, after reviewing the status of the high energy run, to decide on the appropriate time before the end of the 1997 cycle for LEP to change to operation at 130/136 GeV. The 4-jet WG is invited to present the criteria at this meeting for ruling out or confirming a signal in the new data. After 6 pb-1 have been collected at low energy, a second meeting should be held to compare the data with the pre-defined criteria and to decide on a continuation at low energy or a return to high energy. This implies the need for the four LEP detectors to carry out their 4-jet analyses in quasi real time.

6. Any other business

The Chairman announced that he had received a letter indicating an interest to develop a proposal for direct measurement of the LEP2 beam energy by Moller scattering. The apparatus involves a gas jet, silicon detectors and an electromagentic calorimeter. The expected error in sigma(Ebeam) is about 2 MeV at the position of the experiment and 7 MeV when corrected to the positions of the detectors.

In view of the unique nature of the LEP experimental data, there is strong interest for them to be archived for possible future use. A new LEP working group is therefore being established to coordinate the archiving of the LEP experimental data and the preparation of the associated documentation. The four LEP collaborations are invited to nominate two members each for this group, which will be formed together with representatives from IT Division.

8. Next LEPC meeting

The dates of the next meeting of the LEPC are Tuesday-Wednesday, 11-12 November 1997. The tentative dates for the LEPC in 1998 are

31 March - 1 April 31 August (with 27-28 August in reserve) 12 - 13 November
 

DOCUMENTS RECEIVED

[1] ALEPH Collaboration, A proposal for a short LEP run at a centre-of-mass energy of 130-136 GeV, LEPC 97-7.